What is the variable for 6(7-6d)+20=-26
Answer by mananth(16949)

6(7-6d)+20=-26
42-36d+20=-26
62-36d=-26
add-62
-62+62-36d=-26-62
-36d=-88
/-36
d=-88/-36
d=22/9
..
m.ananth@hotmail.ca